Determine the convergence or divergence of the sequence. (I know it diverges, but I need a detailed steps and explanation for this.)

n 31. an = = (-1) n + 1
Answer #1

(31) Given an an = (-1) (m 44 Thus L = lin an= Lim (-1)^(-7 n700 no no1 = Lin (1) x nas *(1+1/n) Lien an afn is even =1 -1